Ball A with mass of 100 g moving in a constant velocity of 200 cm /s, hits Ball B with a mass of 400 g which is at rest. As a result of the collision, Ball A bounced back at 120 cm/s. determine the final velocity of Ball B.

1
Expert's answer
2021-11-14T17:14:35-0500

We can find the final velocity of ball B from the law of conservation of momentum:


"m_Av_{Ai}=-m_Av_{Af}+m_Bv_{Bf},""v_{Bf}=\dfrac{m_A(v_{Ai}+v_{Af})}{m_B},""v_{Bf}=\dfrac{100\ g\times(200\ \dfrac{cm}{s}+120\ \dfrac{cm}{s})}{400\ g}=80\ \dfrac{cm}{s}."